OpenStudy (nwolf):

You are getting ready to purchase your first home. The sales representative here tells you they have two floor plans available, and they have 40 homes still for sale. The equation for this is x + y = 40. The representative also tells you that floor plan one sells for $185,000 and floor plan two sells for $220,000. She also mentions that all the available houses combined are worth $8,170,000. The equation for this is185000x + 220000y = 8,170,000

OpenStudy (anonymous):

Is it the entire post?

OpenStudy (nwolf):

You are getting ready to purchase your first home. The sales representative here tells you they have two floor plans available, and they have 40 homes still for sale. The equation for this is x + y = 40. The representative also tells you that floor plan one sells for $185,000 and floor plan two sells for $220,000. She also mentions that all the available houses combined are worth $8,170,000. The equation for this is185000x + 220000y = 8,170,000

OpenStudy (nwolf):

Use elimination to determine how many houses are available in each floor plan using the given equations.

OpenStudy (anonymous):

x + y = 40 185x + 220y = 8,170 Multiply 220 to the first equation: 220x + 220 y = 8,800 Subtract with the second equa. -> 35x = 630 => x = 630/ 35 = 18 houses on on first floor. Thus y = 40 - 18 = 22 houses on the second floor